OpenStudy (anonymous):

Consider the differential equation dy/dt = -2ty^2. Calculate its general solution and find all values of Yo such that the solution to the intial value problem y(-1)=Yo does not blow up or down in finite time. In other words, find all Yo such that the solution is defined for all real t. @dumbcow @Nurali @zepdrix @Luigi0210 @Compassionate @whpalmer4 @ajprincess @tkhunny @e.mccormick @hba @nincompoop @Zale101 @SithsAndGiggles @Ashleyisakitty @rose21

OpenStudy (nincompoop):

what is Yo?

OpenStudy (nincompoop):

Y naut?

OpenStudy (anonymous):

Yeah

OpenStudy (dumbcow):

you need to separate variables ... get y terms with dy and t terms with dt then integrate both sides

OpenStudy (anonymous):

Ok I got the general solution y(t) = 1/(1+c)..now can you please tell me how to solve the problem.

OpenStudy (nincompoop):

how did you get that?

OpenStudy (nincompoop):

int by substitution

OpenStudy (dumbcow):

you need a variable in the solution ?

OpenStudy (anonymous):

y(t) = 1/(t^2+c)

OpenStudy (dumbcow):

ok now plug in -1 for t \[Y_0 = \frac{1}{1+C}\] oh i see what you posted earlier

OpenStudy (anonymous):

The second part was the main part that is really making me mad..please help

OpenStudy (dumbcow):

solve for C then see when Y0 makes C undefined

OpenStudy (anonymous):

C is never undefined because it's arbitrary..lmao

OpenStudy (nincompoop):

solve for c so that y(t) >= 0

OpenStudy (dumbcow):

haha roflmao :) anyway \[C = \frac{1}{Y_0} -1\] Y0 cant be 0

OpenStudy (nincompoop):

I wish I can medal you again @dumbcow

OpenStudy (anonymous):

Why must y(t) be greater than zero? Why can't it be less than zero?

OpenStudy (nincompoop):

negative time?

OpenStudy (anonymous):

We aren't talking about time here we're talking about the solutions.

zepdrix (zepdrix):

\[\Large\bf\sf C=\frac{y_o-1}{y_o}\] Hmm I'm thinking Yo also can't be 1, right? Yo=1 leads to C=0. Which causes a problem for our function right at the start! It can't even get started, because t=0 is undefined.\[\Large\bf\sf y=\frac{1}{0+t^2}\]

OpenStudy (anonymous):

Am I supposed to look at the general solution y(t) = 1/(t^2+c) to notice that c can't be negative otherwise there would be a discontinuity at some value of time?

OpenStudy (nincompoop):

maybe there's a typo

OpenStudy (anonymous):

Then use this finding to my particular solution y(-1) = 1/(1+c) and know that c>0 to find out that the domain of my particular solution is (0,1)?????????

zepdrix (zepdrix):

Ya I think that's a good thing to notice! Any negative c is giving you a discontinuity at some positive t.

OpenStudy (anonymous):

DOes that mean since c>0 that my solution y(-1) can only have Yo values between 1/1 and 1/oo?

OpenStudy (anonymous):